OpenStudy (anonymous):

One number is two more than a second number. Twice the first is 4 less than 3 times the second. Find the numbers.

OpenStudy (e.mccormick):

You have two unknowns in two equations. If you write out the equations, then any valid method for solving for two unknowns will work.

OpenStudy (***[isuru]***):

hi, if the first number is "x" the second number is " x + 2" and twice the first is " 2x" also... 3 less than 3 times the second no. is " 2(x+2) -3" so... the problem says Twice the first is 4 less than 3 times the second. which means.. 2x = 3(x+2) -4 now solve this for "x" wt will u get ?

OpenStudy (austinl):

x+2=y 2x+4=3y I believe this is what is being asked.... A system of equations?

OpenStudy (anonymous):

\[\{x=2+y,2x=3y-4\}\]

OpenStudy (anonymous):

@austinL Yes, I need a system of equations. Let me try that.

OpenStudy (anonymous):

{x = 10, y = 8}

OpenStudy (anonymous):

@austinL I just solved it, and it's not quite working. @robtobey yours is working perfectly. Thank you!

OpenStudy (***[isuru]***):

@austinL ummmm... ur second equation seems to have some error bro... the question says.. twice the first number which is "2y" according to ur system of equation is equal to 4 less than thrice the second number... so.... it should be... 2y = 4 - 3(x+2)

OpenStudy (anonymous):

Liv_16 Thank you for the medal.

OpenStudy (anonymous):

@robtobey Thank YOU for the help =)

OpenStudy (e.mccormick):

@***[ISURU]*** \(2x+4=3y \implies 2x=3y-4\) Both mean "Twice the first is 4 less than 3 times the second. "

OpenStudy (austinl):

Lol, mine is just subtly different than rob's. I put "+4" on the left, and he put "-4" on the right. Though, subtle differences make big changes in mathematics. +1 for you @robtobey @***[ISURU]*** What you have just done is the next step in solving the system of equations. Substitution.